OpenStudy (anonymous):

(1-1/2) (1-1/3) (1-1/4).........(1-1/100) is =? A.1/120, b.1/108 c.1/100. d.1/50.

OpenStudy (anonymous):

c. what is the product of the 1st and 2nd term? (1-1/2)*(1-1/3) = .333333 what is the product of the first 3 terms (1-1/2) (1-1/3) (1-1/4) = .25 what pattern do you see here? the answer to the product of the terms is 1- the last bracket. as in the examples above 1- (1-1/3) =.33333 1-(1-1/4) = .25 so to multiply all the way to 1/100 would be 1-(1-1/100) = answer